OpenStudy (anonymous):

(x+2)square=Xsquare-6x+1 Slove the equation algebraically

myininaya (myininaya):

\[(x+2)^2=x^2+4x+4\] --------------------- \[x^2+4x+4=x^2-6x+1\] subtract x^2 on both sides \[4x+4=-6x+1\] you got it now?